# GSPRec: Temporal-Aware Graph Spectral Filtering for Recommendation

### v1

@@ 3.3 Spectral Filtering Framework
To model user preferences at different levels of specificity, we perform spectral filtering on the Laplacian graph $\mathbf{L}=\mathbf{U}\boldsymbol{\Lambda}\mathbf{U}^{T}$ where $\boldsymbol{\Lambda}=\text{diag}(\lambda_{1},\ldots,\lambda_{m+n})$ contains eigenvalues in ascending order $0\leq\lambda_{1}\leq\lambda_{2}\leq\ldots\leq\lambda_{m+n}\leq 2$ and $\mathbf{U}=[\mathbf{u}_{1},\ldots,\mathbf{u}_{m+n}]$ contains the corresponding eigenvectors. In practice, for computational efficiency, we perform a truncated eigendecomposition focusing on $r\ll(m+n)$ which includes the low- and mid-frequency bands relevant to our filters. The lowpass filter $g_{\text{LP}}(\lambda)$ and bandpass filter $g_{\text{BP}}(\lambda)$ operate on the corresponding $r$ eigenpairs by selecting and amplifying the mid-frequency signals present within this subspace.
@@
To capture mid-frequency user-specific behavior, we design a bandpass filter to the graph spectrum using a Gaussian kernel:
@@
where $c$ and $w$ control the center and width of the frequency band. This form is motivated by our empirical analysis (Appendix G), which indicates that spectral coefficients for intermediate eigenvalues often show higher magnitude for user-specific preferences compared to global trends. Let $\mathbf{G}_{\text{BP}}=\text{diag}(g_{\text{BP}}(\lambda_{1}),\ldots,g_{\text{BP}}(\lambda_{r}))$ denote the spectral bandpass filter. We filter the item-degree-normalized user-item interaction signal using the band-pass operator:
@@
In our framework, mid-frequency components correspond to more user-specific variations, while low frequencies capture smoother, population-level preferences. We compute item scores by fusing mid-frequency personalization with low-frequency global trends. Specifically, we compute:
@@ 4.1 Setup
We conduct experiments on four benchmark datasets spanning two domains: (1) ML100K, ML1M [13] and Netflix [1] from the movie recommendation domain; (2) Amazon Beauty [32] from the e-commerce product domain. These datasets were selected to maintain comparability with recent spectral methods [52, 53] that established them as standard benchmarks for this class of approaches. All datasets are divided into training set and test set with the ratio of 8:2, and we take 10% of the training set as validation set for hyperparameter tuning. Table 3 shows the statistics of these datasets.
@@
We compare GSPRec with: (i) Popularity; (ii) GCN-based methods: LightGCN [14], LR-GCCF [4], IMP-GCN [23], SimpleX [28], UltraGCN [29]; and (iii) GSP-based methods: GF-CF [42], PGSP [25], FaGSP [52], HiGSP [53]. A detailed description of these baselines is provided in Appendix E. Unlike existing GSP methods that use only collaborative signals and low-pass filtering, GSPRec encodes temporal transitions into the graph topology and employs a dual-filter approach targeting both low and bandpass filter components.
@@
Metrics.
Following prior work [14, 42, 52, 53, 25], we use Normalized Discounted Cumulative Gain (NDCG@$k$) and Mean Reciprocal Rank (MRR@$k$) at $k\in\{5,10,20\}$. We apply spectral filtering to obtain preference scores, exclude previously interacted items, and rank the top-$k$ items.
4.2 Overall Performance Comparison
Table 4: Performance comparison on four public datasets. The best performance is denoted in bold. “M@K” and “N@K” are short for “MRR@K” and “NDCG@K”, respectively. “Improv.” denotes the percentage improvement of GSPRec compared to the strongest baseline method.
@@ 4.2 Overall Performance Comparison
Popularity LightGCN LR-GCCF IMP-GCN SimpleX UltraGCN GF-CF PGSP FaGSP HiGSP GSPRec
ML100K N@5 0.5672 0.7034 0.5587 0.6764 0.6995 0.6786 0.6875 0.6851 0.7106 0.7166 0.7543 +5.26%
M@5 0.5151 0.6362 0.4943 0.6006 0.6397 0.6094 0.6234 0.6149 0.6518 0.6629 0.7358 +11.00%
N@10 0.5857 0.6771 0.5603 0.6605 0.6866 0.6688 0.6843 0.6722 0.7092 0.7021 0.7572 +6.77%
M@10 0.5296 0.5877 0.4616 0.5690 0.6064 0.5749 0.6010 0.5767 0.6384 0.6380 0.7443 +16.59%
N@20 0.5867 0.6618 0.5555 0.6605 0.6543 0.6486 0.6621 0.6495 0.6989 0.6724 0.7465 +6.81%
M@20 0.5315 0.5652 0.4411 0.5690 0.5411 0.5289 0.5593 0.5496 0.5901 0.5782 0.7457 +26.37%
ML1M N@5 0.0958 0.5845 0.3540 0.5583 0.5934 0.5739 0.5935 0.5963 0.6112 0.6062 0.6237 +2.05%
M@5 0.1943 0.5160 0.2997 0.4934 0.5244 0.5087 0.5254 0.5313 0.5430 0.5386 0.5715 +5.25%
N@10 0.0960 0.5873 0.3787 0.5594 0.5921 0.5773 0.5897 0.5923 0.6082 0.6042 0.6431 +5.74%
M@10 0.2180 0.5010 0.2946 0.4685 0.5051 0.4886 0.4996 0.5063 0.5229 0.5148 0.5837 +11.63%
N@20 0.1035 0.5692 0.3980 0.5416 0.5687 0.5600 0.5678 0.5687 0.5808 0.5769 0.6431 +10.73%
M@20 0.2309 0.4644 0.2924 0.4292 0.4625 0.4500 0.4557 0.4646 0.4748 0.4620 0.5871 +23.65%
Netflix N@5 0.5105 0.6822 0.6023 0.6328 0.6603 0.5680 0.7005 0.6811 0.7162 0.7162 0.7201 +0.54%
M@5 0.4505 0.6147 0.5323 0.5528 0.5902 0.4941 0.6360 0.6127 0.6524 0.6532 0.6680 +2.27%
N@10 0.5488 0.6814 0.6134 0.6307 0.6694 0.5746 0.6928 0.6756 0.7079 0.7062 0.7270 +2.70%
M@10 0.4712 0.5974 0.5234 0.5386 0.5839 0.4708 0.6126 0.5905 0.6293 0.6275 0.6762 +7.45%
N@20 0.5604 0.6642 0.6043 0.6152 0.6572 0.5580 0.6689 0.6543 0.6814 0.6800 0.7213 +5.86%
M@20 0.4773 0.5636 0.4963 0.4995 0.5555 0.4309 0.5691 0.5502 0.5827 0.5800 0.6777 +16.30%
Beauty N@5 0.0149 0.0668 0.0533 0.0570 0.0682 0.0599 0.0659 0.0674 0.0712 0.0718 0.0733 +2.09%
M@5 0.0118 0.0530 0.0423 0.0444 0.0544 0.0476 0.0523 0.0538 0.0564 0.0574 0.0619 +7.84%
N@10 0.0200 0.0763 0.0635 0.0667 0.0761 0.0649 0.0751 0.0765 0.0799 0.0771 0.0879 +10.01%
M@10 0.0139 0.0519 0.0443 0.0451 0.0517 0.0441 0.0517 0.0525 0.0549 0.0531 0.0680 +23.86%
N@20 0.0256 0.0829 0.0678 0.0729 0.0819 0.0689 0.0819 0.0828 0.0846 0.0799 0.1013 +19.74%
M@20 0.0154 0.0466 0.0388 0.0409 0.0461 0.0393 0.0469 0.0481 0.0485 0.0456 0.0718 +48.04%
@@
Table 4 presents performance comparisons. GSPRec consistently outperforms all baselines across all metrics and datasets. Three Key observations: (1) consistent NDCG@$k$/MRR@$k$ improvements show stronger ranking quality, (2) larger gains on sparse datasets confirm robustness, and (3) improvements over recent GSP-based methods validate the benefit of leveraging both sequential and spectral signals.
4.3 Parameter Study
Sequential-Aware Graph:
To isolate the contribution of sequential graph, we compare our full model with four variants shown in Table 6: (1) GSPRec-NB: no band-pass filter, (2) GSPRec-NL: no low-pass filter, (3) GSPRec-NS: no sequential diffusion, (4) GSPRec-SE: sequence edges only (no diffusion). The non-sequential variant (GSPRec-NS) shows significant performance drops, confirming the value of sequential information. The direct-sequence variant (GSPRec-SE) performs slightly worse than the full model, demonstrating that higher-order transitions captured by multi-hop diffusion provide additional benefits.
@@ 4.3 Parameter Study
$\mathbf{F}_{\text{LP}}$+$\mathbf{F}_{\text{BP}}$ ($\phi$=0.3) 0.6431 0.0074 0.5837 0.0081
$\mathbf{F}_{\text{LP}}$+$\mathbf{F}_{\text{BP}}$ ($\phi$=0.5) 0.6327 0.0076 0.5714 0.0085
$\mathbf{F}_{\text{LP}}$+$\mathbf{F}_{\text{BP}}$ ($\phi$=0.7) 0.6281 0.0078 0.5682 0.0087
Spectral Filtering Performance:
Table 3 shows optimal filter settings across datasets: (1) larger datasets require higher eigendecomposition dimensions $r$; (2) sparser datasets benefit from wider bandpass filters; and (3) fusion weight $\phi$ correlates with dataset density. Figure 4 illustrates filter frequency responses spanning both global and user-specific bands. As shown in Table 6, the dual-filter approach consistently outperforms single-filter alternatives, with best performance at $\phi$=0.3 on ML1M. Notably, the bandpass filter alone outperforms the lowpass filter alone by 4.7% in N@10 and 1.6% in M@10, highlighting the importance of mid-frequency components for personalization. Ablation results in Table 6 further support this, with GSPRec-NB (no band-pass) showing the largest performance drops.
4.4 Parameter Sensitivity Analysis
@@
Figures 5 and 6 show how performance varies with band-pass filter parameters. The center position ($c$) exhibits dataset-specific optima, reflecting varying spectral distributions across recommendation domains. Filter width ($w$) interacts with center position. The fusion weight $\phi$ also shows dataset-dependent patterns: denser datasets favor global signals (lower $\phi$), while sparser datasets benefit from stronger mid-frequency personalization (higher $\phi$).
@@ 5 Conclusion
We introduced GSPRec, a graph–spectral framework that jointly models collaborative and sequential signals via a unified Laplacian. Our method augments the user–item graph with multi-hop, exponentially decayed item transitions and applies dual spectral filters. Theoretically, we ensure spectral convergence and Laplacian validity. Empirically, we demonstrate consistent improvements across four benchmarks over state-of-the-art GSP methods. Ablation studies confirm the necessity of both sequential graph encoding and frequency-aware filtering. Future work includes extending to directed graphs, learning adaptive spectral filters, and applying to diverse recommendation contexts.
6 Broader Impact
This work improves recommender systems by integrating sequential behavior with graph spectral analysis, benefiting users across e-commerce, media, and education domains. However, spectral methods may amplify biases in historical data, potentially reinforcing popularity bias or increasing computational overhead. Future work should explore fairness-aware spectral filters, robustness evaluations, and privacy-preserving techniques. We encourage practitioners to consider these tradeoffs when deploying these systems in real-world settings.
@@ 6 Broader Impact
This supplementary material provides detailed information that complements our main paper. It includes formal notation definitions, theoretical analysis, algorithmic details, computational complexity analysis, and experimental information to support reproducibility.
@@ Appendix A Table of Notations
$\mathbf{U},\boldsymbol{\Lambda}$ Matrices of eigenvectors and eigenvalues of $\mathbf{L}$, respectively
@@
$\mathbf{F}_{\text{LP}}$ Lowpass filtered user-item signal

### v2

@@ 3.3 Spectral Filtering Framework
To model user preferences at different levels of specificity, we perform spectral filtering on the Laplacian graph $\mathbf{L}=\mathbf{U}\bm{\Lambda}\mathbf{U}^{T}$ where $\bm{\Lambda}=\text{diag}(\lambda_{1},\ldots,\lambda_{m+n})$ contains eigenvalues in ascending order $0\leq\lambda_{1}\leq\lambda_{2}\leq\ldots\leq\lambda_{m+n}\leq 2$ and $\mathbf{U}=[\mathbf{u}_{1},\ldots,\mathbf{u}_{m+n}]$ contains the corresponding eigenvectors. In practice, for computational efficiency, we perform a truncated eigendecomposition focusing on $r\ll(m+n)$ which includes the low- and mid-frequency bands relevant to our filters. The lowpass filter $g_{\text{LP}}(\lambda)$ and bandpass filter $g_{\text{BP}}(\lambda)$ operate on the corresponding $r$ eigenpairs by selecting and amplifying the mid-frequency signals present within this subspace.
@@
To capture mid-frequency user-specific behavior, we design a bandpass filter using a Gaussian kernel:
@@
where $c$ and $w$ control the center and width of the frequency band. Our empirical analysis (Appendix G) shows spectral coefficients for intermediate eigenvalues have higher magnitude for user-specific preferences than global trends. This aligns with spectral theory: low frequencies capture global popularity (smooth across graph), high frequencies represent noise (rapid variation), while mid-frequencies are piecewise smooth, which varies between communities but consistent within them. In recommendations, user communities correspond to groups with shared item preferences, making mid-frequency the natural band for personalization.
Let $\mathbf{G}_{\text{BP}}=\text{diag}(g_{\text{BP}}(\lambda_{1}),\ldots,g_{\text{BP}}(\lambda_{r}))$ denote the spectral bandpass filter. We filter the item-degree-normalized user-item interaction signal:
@@
In our framework, mid-frequency components correspond to more personalized patterns, while low frequencies capture smoother, population-level preferences. We compute item scores by fusing mid-frequency personalization with low-frequency global trends. Specifically, we compute:
@@
Both $\mathbf{F}_{\text{BP}}$ and $\mathbf{F}_{\text{LP}}$ are outputs of linear spectral filters applied to the graph Laplacian $\mathbf{L}$. Their convex combination remains a valid linear operator, as the space of spectral filters is closed under addition and scalar multiplication Sandryhaila & Moura (2013); Püschel & Moura (2006). Figure 2 illustrates our dual-filter architecture.
@@ 4.1 Experimental Setup
We conduct experiments on four benchmark datasets: ML100K, ML1M Harper & Konstan (2015), Netflix Bennett & Lanning (2007), and Amazon Beauty Ni et al. (2019). These datasets maintain comparability with recent spectral methods Xia et al. (2025; 2024). We use an 8:1:1 train/test split with 10% of training data for validation. Statistics are shown in Table 3.
@@
We compare GSPRec with: (i) Popularity; (ii) GCN-based methods: LightGCN He et al. (2020), LR-GCCF Chen et al. (2020b), IMP-GCN Liu et al. (2021), SimpleX Mao et al. (2021a), UltraGCN Mao et al. (2021b); and (iii) GSP-based methods: GF-CF Shen et al. (2021), JGCF Guo et al. (2023), PGSP Liu et al. (2023), HiGSP Xia et al. (2024), FaGSP Xia et al. (2025). Detailed descriptions are in Appendix E. GSPRec leverages temporal graph construction with dual-filter approach targeting both low- and band-pass components.
@@
Metrics:
Following prior work He et al. (2020); Shen et al. (2021); Xia et al. (2025); Guo et al. (2023); Xia et al. (2024); Liu et al. (2023), we use Normalized Discounted Cumulative Gain (NDCG@$k$) and Mean Reciprocal Rank (MRR@$k$) at $k\in\{5,10,20\}$, abbreviated as N@k and M@k respectively. We apply spectral filtering to obtain preference scores, exclude previously interacted items, and rank the top-$k$ items.
4.2 Overall Performance
Table 4: Performance comparison on four public datasets. The best performance is denoted in bold. “Improv.” denotes the percentage improvement of GSPRec compared to the strongest baseline method.
@@ 4.2 Overall Performance
Popularity LightGCN LR-GCCF IMP-GCN SimpleX UltraGCN GF-CF JGCF PGSP FaGSP HiGSP GSPRec
ML100K N@5 0.5672 0.7034 0.5587 0.6764 0.6995 0.6786 0.6875 0.6582 0.6851 0.7106 0.7166 0.7543 +5.26%
M@5 0.5151 0.6362 0.4943 0.6006 0.6397 0.6094 0.6234 0.4471 0.6149 0.6518 0.6629 0.7358 +11.00%
N@10 0.5857 0.6771 0.5603 0.6605 0.6866 0.6688 0.6843 0.6695 0.6722 0.7092 0.7021 0.7572 +6.77%
M@10 0.5296 0.5877 0.4616 0.5690 0.6064 0.5749 0.6010 0.4222 0.5767 0.6384 0.6380 0.7443 +16.59%
N@20 0.5867 0.6618 0.5555 0.6605 0.6543 0.6486 0.6621 0.6730 0.6495 0.6989 0.6724 0.7465 +6.81%
M@20 0.5315 0.5652 0.4411 0.5690 0.5411 0.5289 0.5593 0.4197 0.5496 0.5901 0.5782 0.7457 +26.37%
ML1M N@5 0.0958 0.5845 0.3540 0.5583 0.5934 0.5739 0.5935 0.6121 0.5963 0.6112 0.6062 0.6237 +2.05%
M@5 0.1943 0.5160 0.2997 0.4934 0.5244 0.5087 0.5254 0.4172 0.5313 0.5430 0.5386 0.5715 +5.25%
N@10 0.0960 0.5873 0.3787 0.5594 0.5921 0.5773 0.5897 0.6238 0.5923 0.6082 0.6042 0.6431 +3.10%
M@10 0.2180 0.5010 0.2946 0.4685 0.5051 0.4886 0.4996 0.3869 0.5063 0.5229 0.5148 0.5837 +11.63%
N@20 0.1035 0.5692 0.3980 0.5416 0.5687 0.5600 0.5678 0.6270 0.5687 0.5808 0.5769 0.6431 +2.57%
M@20 0.2309 0.4644 0.2924 0.4292 0.4625 0.4500 0.4557 0.3730 0.4646 0.4748 0.4620 0.5871 +23.65%
Netflix N@5 0.5105 0.6822 0.6023 0.6328 0.6603 0.5680 0.7005 0.7065 0.6811 0.7162 0.7162 0.7201 +0.54%
M@5 0.4505 0.6147 0.5323 0.5528 0.5902 0.4941 0.6360 0.5283 0.6127 0.6524 0.6532 0.6680 +2.27%
N@10 0.5488 0.6814 0.6134 0.6307 0.6694 0.5746 0.6928 0.7148 0.6756 0.7079 0.7062 0.7270 +1.71%
M@10 0.4712 0.5974 0.5234 0.5386 0.5839 0.4708 0.6126 0.4837 0.5905 0.6293 0.6275 0.6762 +7.45%
N@20 0.5604 0.6642 0.6043 0.6152 0.6572 0.5580 0.6689 0.7164 0.6543 0.6814 0.6800 0.7213 +0.68%
M@20 0.4773 0.5636 0.4963 0.4995 0.5555 0.4309 0.5691 0.4273 0.5502 0.5827 0.5800 0.6777 +16.30%
Beauty N@5 0.0149 0.0668 0.0533 0.0570 0.0682 0.0599 0.0659 0.0501 0.0674 0.0712 0.0718 0.0733 +2.09%
M@5 0.0118 0.0530 0.0423 0.0444 0.0544 0.0476 0.0523 0.0535 0.0538 0.0564 0.0574 0.0619 +7.84%
N@10 0.0200 0.0763 0.0635 0.0667 0.0761 0.0649 0.0751 0.0612 0.0765 0.0799 0.0771 0.0879 +10.01%
M@10 0.0139 0.0519 0.0443 0.0451 0.0517 0.0441 0.0517 0.0590 0.0525 0.0549 0.0531 0.0680 +15.25%
N@20 0.0256 0.0829 0.0678 0.0729 0.0819 0.0689 0.0819 0.0721 0.0828 0.0846 0.0799 0.1013 +19.74%
M@20 0.0154 0.0466 0.0388 0.0409 0.0461 0.0393 0.0469 0.0622 0.0481 0.0485 0.0456 0.0718 +15.43%
@@
Table 4 presents performance comparisons. GSPRec consistently outperforms all baselines across metrics and datasets, with improvements over recent GSP-based methods validating the effectiveness of leveraging sequential graph construction with mid-frequency spectral filtering.
Performance improvements range from 0.54% to 26.37% across datasets and metrics. Beauty dataset achieves 19.74% improvement on N@20, while ML100K shows improvements ranging from 5.26% to 26.37% and Netflix demonstrates gains of 0.54% to 16.30%. This variation indicates that method effectiveness depends on dataset-specific characteristics that extend beyond simple density metrics. The performance variation correlates with dataset-specific optimal parameter configurations shown at Table 3. ML100K exhibits optimal performance at lower frequency centers, Netflix at mid-range frequencies, while Beauty and ML1M favor higher frequencies. These indicate that GSPRec adapts to the underlying spectral structure of each dataset rather than imposing a fixed frequency range.
4.3 Ablation Analysis
@@
$\mathbf{F}_{\text{LP}}$$\mathbf{F}_{\text{BP}}$ ($\phi$=0.3) 0.6431 0.0074 0.5837 0.0081
$\mathbf{F}_{\text{LP}}$$\mathbf{F}_{\text{BP}}$ ($\phi$=0.5) 0.6327 0.0076 0.5714 0.0085
$\mathbf{F}_{\text{LP}}$$\mathbf{F}_{\text{BP}}$ ($\phi$=0.7) 0.6281 0.0078 0.5682 0.0087
Table 6 reveals a clear component hierarchy. Removing bandpass filtering causes the largest performance degradation, followed by lowpass filtering removal, while sequential diffusion removal causes the smallest drop. This hierarchy demonstrates that spectral filtering strategy contributes more substantially to performance than graph augmentation through sequential information. Table 6 confirms that bandpass filtering alone outperforms lowpass alone by 4.7%, supporting the effectiveness of mid-frequency targeting over traditional low-frequency approaches.
Parameter analysis reveals both universal architectural constants and dataset-specific adaptations. Diffusion parameters $\alpha=0.4$ and $d=2$ remain across all datasets, which suggests fundamental properties of sequential transition structures that transcend domain-specific characteristics. However, spectral parameters exhibit systematic variation that eigenspace dimensionality $r$ scales linearly with dataset size, while bandpass center positions cluster according to dataset characteristics, in which indicating that different datasets exhibit distinct spectral signatures in their interaction patterns.
The fusion weight $\phi$ provides interpretable insight into the balance between personalization and popularity signals. The optimal value of 0.3 on ML1M indicates that bandpass-based personalization dominates the recommendation process while lowpass popularity signals provide complementary information. This quantifies the relative importance of different spectral components and demonstrates how GSPRec achieves effective signal separation.
4.4 Parameter Sensitivity
@@
Figures 4 and 5 demonstrate that optimal filter configurations depend on both center position and dataset characteristics. The interaction between center and width parameters confirms that different datasets require distinct spectral windows to capture personalized patterns while suppressing noise amplification. This adaptive behavior explains why fixed low-pass or high-pass approaches fail to achieve consistent improvements across diverse recommendation scenarios.
Notably, the sensitivity curves exhibit relative flatness around optimal parameter values, indicating that moderate parameter deviations do not substantially degrade performance. This robustness suggests that GSPRec does not require precise hyperparameter tuning.
4.5 Computational Analysis
Table 7 shows that GSPRec requires 1.27 minutes on ML1M, comparable to other GSP methods (1.18-2.36 minutes) and substantially more efficient than GCN approaches (16-135 minutes). This positions GSPRec on the accuracy-efficiency frontier, unlike GCN methods that achieve moderate improvements at high computational cost, or GSP methods that sacrifice accuracy for speed, GSPRec provides both computational efficiency and high accuracy by targeting the most informative bands.
Table 7: Runtime comparison on ML1M dataset (minutes).
Method LightGCN SimpleX UltraGCN GF-CF PGSP FaGSP HiGSP GSPRec
Runtime 135.96 109.83 16.09 1.18 2.36 2.08 2.04 1.27
@@ 5 Conclusion
We presented GSPRec, a temporal-spectral collaborative filtering framework that integrates sequential transitions with dual-frequency spectral filtering. our approach consistently outperforms baselines across four datasets by targeting mid-frequency components, with bandpass filtering achieving 4.7% improvement over state-of-the-art methods.
Limitations and Future Directions: The symmetric construction trades directional information for spectral stability, and eigendecomposition creates computational bottlenecks addressable through polynomial filtering. This work demonstrates that restructuring graph spectra through sequential encoding enables more effective personalization than filtering static user-item graphs.
@@ Appendix A Table of Notations
$\mathbf{U},\bm{\Lambda}$ Matrices of eigenvectors and eigenvalues of $\mathbf{L}$, respectively
@@
$\mathbf{F}_{\text{LP}}$ Low-pass filtered user-item signal